Election of Council Delegates

As announced in the previous issue of the Newsletter, nominations are required for Council delegates representing individual members of the Society. On 1 November 1995 there were 1674 individual members, and this means that there should be 17 delegates representing them. During the term ending with this year the delegates representing individual members were:

Term 1992–95: Aubin, Thierry; Bérard, Pierre; Branner, Bodil; Deshouillers, Jean-Marc; Desolneux-Moulis, N; Dineen, Sean; Fenn, Roger Andrew; Howie, John M; Karoubi, Max; Malliavin, Paul; Martin-Deschamps, Mireille.

Term 1994-97: Castellet, Manuel; Jaiani, George; Kufner, Alois; Langevin, Remi; Michor, Peter W; Schwarz, Wolfgang.

The terms of the first mentioned 11 delegates ends now. They are all eligible for re-election but Professors Deschamps, Howie and Malliavin have informed the Secretary that they do not wish to continue as delegates. Nominations are now sought for 11 delegates to serve for the years 1996–1999. Attached to this notice is a nomination form. Completed nomination forms must arrive at the Society’s office in Helsinki by 10 February 1996. If there are more nominations than the allowed number of delegates, a postal election will be held; members will receive ballot forms at the end of February 1996 and these must be returned by 10 April 1996. Nominated delegates must be individual members of the Society and they must be proposed and seconded by individual members. The Society will pay subsistence costs for them to attend the Council meetings, if needed, but is not able to cover travel costs except perhaps in cases of particular hardship. Candidates for election are invited to submit with their nomination form a short biography (not more than 200 words) together with a Statement of not more than 100 words in support of their candidature. These will be circulated to the members with the ballot forms. A copy of these can be sent as a text file by e-mail to the Secretariat to makelainen@cc.helsinki.fi.

Peter Michor, Secretary

Nomination Form for Council Delegate

NAME: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

TITLE: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

ADDRESS: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .
. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

PROPOSER: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

SECONDER: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . I certify that I am an individual member of the EMS and that I am willing to stand for election as a delegate of individual members to the Council. SIGNATURE OF CANDIDATE:

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

DATE: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .

Completed forms should be sent to:

Ms T. Mäkeläinen EMS Secretariat Department of Mathematics P.O. Box 4 FIN-00014 University of Helsinki Finland to arrive by February 10, 1996. A photocopy of this form is acceptable.